# FAIRGAME: a Framework for AI Agents Bias Recognition using Game Theory

### v4

@@ 2.1 Creation and execution of games
Algorithm 1 processes the configuration file CF and the set of prompt templates $PT$ as inputs, and outputs a list $G$ containing all instantiated games. The process involves validating inputs, extracting relevant game information, configuring agents, and creating the games.
@@ 3.2 Games
To explore strategic variations in adversarial interactions, and assess the sensitivity of each LLM to game parameters, we define multiple configurations of the Prisoner’s Dilemma game. Using an established scaling of dilemma strength [54], fixing other payoff matrix entries, the dilemma strength in the Prisoner’s Dilemma decrease with the difference between mutual reward and mutual punishment. For the conventional configuration with penalties $x_{1,1}=(6,6)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, this difference is 2(6)=4-2-(-6)=4. The harsh configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(5,5)$ and the mild configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, have differences of $3$ and $6$, respectively.
@@ 4.1 Prisoner’s dilemma
In most cases, the preferred end result favors agents defecting, as suggested by game theory. This aligns with the Nash equilibrium of the Prisoner’s Dilemma, where mutual defection is the dominant strategy. Nevertheless, there are inconsistencies across languages and combinations of personalities, which suggests that the agents’ behavior is influenced by factors beyond the payoff matrix, such as languages and inherent biases present in LLM training data. For instance, penalties are generally lower in English, particularly for GPT-4o and Claude 3.5 Sonnet, when the number of rounds is unknown, indicating more consistent cooperative behavior in their primary training language. Broader variability in languages like French or Arabic with high penalties suggests challenges in interpreting the game dynamics due to linguistic or cultural differences, while penalties remain high in Mandarin Chinese and Vietnamese, particularly for Mistral Large. In mixed personality settings (CS), selfish agents exploit cooperative ones, leading to higher penalties for the latter, consistent with game-theoretic predictions. For selfish pairings (SS), penalties are high but exhibit low variability, as mutual defection is the rational choice. When the number of rounds is unknown, penalties are higher across all settings, reflecting uncertainty that discourages cooperation. Conversely, when rounds are known, penalties decrease, particularly in CC and CS settings, as agents can plan strategies with the endgame in mind. Claude 3.5 Sonnet and GPT-4o show significant reductions in penalties when rounds are known, especially in English, demonstrating their ability to adapt to game structure. In contrast, Mistral Large shows less sensitivity to this information. Overall, knowing the number of rounds promotes cooperation in CC and CS settings, while SS settings remain unaffected, as defection remains the dominant strategy.
Statistically, we recognize that, when ’selfish’ agents are present, there is lower variability in the results, and that some LLMs end up with a broader range of possible outputs than others, also depending on the language used. For instance, all LLM have very narrow distributions in English SS, if the agents know the opponent’s personality, but have larger distributions in French, for the same settings. Also, Mistral Large has reduced variability if agents’ personalities are known, while GPT-4o has overall larger variability. We will quantitatively measure this variability in Sec. 4.3.

### v5

@@ 2.1 Creation and execution of games
Algorithm 1 processes the configuration file CF and the set of prompt templates PTPT as inputs, and outputs a list $G$ containing all instantiated games. The process involves validating inputs, extracting relevant game information, configuring agents, and creating the games.
@@ 3.2 Games
To explore strategic variations in adversarial interactions, and assess the sensitivity of each LLM to game parameters, we define multiple configurations of the Prisoner’s Dilemma game. Using an established scaling of dilemma strength [54], fixing other payoff matrix entries, the dilemma strength in the Prisoner’s Dilemma decrease with the difference between mutual reward and mutual punishment. For the conventional configuration with penalties $x_{1,1}=(6,6)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, this difference is 2(6)=4-2-(-6)=4. The harsh configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(5,5)$ and the mild configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, have differences of 33 and 66, respectively.
@@ 4.1 Prisoner’s dilemma
In most cases, the preferred end result favors agents defecting, as suggested by game theory. This aligns with the Nash equilibrium of the Prisoner’s Dilemma, where mutual defection is the dominant strategy. Nevertheless, there are inconsistencies across languages and combinations of personalities, which suggests that the agents’ behavior is influenced by factors beyond the payoff matrix, such as languages and inherent biases present in LLM training data. For instance, penalties are generally lower in English, particularly for GPT-4o and Claude 3.5 Sonnet, when the number of rounds is unknown, indicating more consistent cooperative behavior in their primary training language. Broader variability in languages like French or Arabic with high penalties suggests challenges in interpreting the game dynamics due to linguistic or cultural differences, while penalties remain high in Mandarin Chinese and Vietnamese, particularly for Mistral Large. In mixed personality settings (CS), selfish agents exploit cooperative ones, leading to higher penalties for the latter, consistent with game-theoretic predictions. For selfish pairings ($S$), penalties are high but exhibit low variability, as mutual defection is the rational choice. When the number of rounds is unknown, penalties are higher across all settings, reflecting uncertainty that discourages cooperation. Conversely, when rounds are known, penalties decrease, particularly in $C$ and CS settings, as agents can plan strategies with the endgame in mind. Claude 3.5 Sonnet and GPT-4o show significant reductions in penalties when rounds are known, especially in English, demonstrating their ability to adapt to game structure. In contrast, Mistral Large shows less sensitivity to this information. Overall, knowing the number of rounds promotes cooperation in $C$ and CS settings, while $S$ settings remain unaffected, as defection remains the dominant strategy.
Statistically, we recognize that, when ’selfish’ agents are present, there is lower variability in the results, and that some LLMs end up with a broader range of possible outputs than others, also depending on the language used. For instance, all LLM have very narrow distributions in English $S$, if the agents know the opponent’s personality, but have larger distributions in French, for the same settings. Also, Mistral Large has reduced variability if agents’ personalities are known, while GPT-4o has overall larger variability. We will quantitatively measure this variability in Sec. 4.3.